OpenStudy (anonymous):

I REALLY NEED SOME HELP GUYS!! 0.25 moles of a hydrocabon Xwith a molar mass of 72 gmol was burnt completely in an excess of oxygen. The volume of CO2 produced was 30 dm3. i)What is the molecular formula of the hydrocarbon x?

OpenStudy (anonymous):

Moles of carbon in CO2=30/22.4 Moles of carbon in 0.25 moles of the compund=30/22.4 Moles of carbon in 1 mole of compund=30*4/22.4.=approx 5 Hence it would be C5H12